OpenStudy (anonymous):

For questions 1-2, apply the quadratic formula to find the roots of the given function, and then graph the function. 1. (6 points) 0 = 3x^2 + 18x + 22 2. (6 points) 0 = -2x^2 + 8x - 3

OpenStudy (anonymous):

1) x= [- 18 +- sqrt(18^2 - 4*3*22) ]/(2*3) x=(-18+-7.7459)/6 2) x= [-8 +- sqrt(8^2-4*-2*-3)]/(2*-2) x= (-8 +- 6.3245)/(-4) that is your answer!

OpenStudy (anonymous):

dont i get 2 ys and 2 xs to graph?

OpenStudy (anonymous):

yes! look at the '+-'. Just make math to get it!

OpenStudy (anonymous):

im sorry i dont understand these ]: or the way you put it.

OpenStudy (anonymous):

first you calculate using '+'. you find first x. then calculate using '-' and you get the second x!

OpenStudy (anonymous):

x1=(-18+7.7459)/6 x2=(-18-7.7459)/6

OpenStudy (anonymous):

um.....ok....

OpenStudy (anonymous):

y =0 in both case because to their x values, the function value is zero. Thats x is called as zeros of function

OpenStudy (anonymous):

but im getting \[x \approx-1.71,-4.29\]